z=(x-y)F(x2+y2) obtain the partial differential equation by eliminating the arbiyrary function from the following equation


1
Expert's answer
2021-09-02T00:19:22-0400

Let us obtain the partial differential equation by eliminating the arbiyrary function from the following equation "z=(x-y)F(x^2+y^2)." Since "z_x'=F(x^2+y^2)+(x-y)F'(x^2+y^2)2x" and "z_y'=-F(x^2+y^2)+(x-y)F'(x^2+y^2)2y," we conclude that "yz_x'=yF(x^2+y^2)+2xy(x-y)F'(x^2+y^2)" and "xz_y'=-xF(x^2+y^2)+2xy(x-y)F'(x^2+y^2)." After substracting we get the equation "yz_x'-xz_y'=yF(x^2+y^2)+xF(x^2+y^2)." Then "yz_x'-xz_y'=(x+y)F(x^2+y^2)," and hence "F(x^2+y^2)=\frac{yz_x'-xz_y'}{x+y}." It follows that "z=(x-y)F(x^2+y^2)=(x-y)\frac{yz_x'-xz_y'}{x+y}." Consequently, the partial differential equation is the following:

"z(x+y)=(x-y)(yz_x'-xz_y')."